# Matrix harmonic analysis at high temperature via the Dirichlet process

### v1

@@ 1.2.5. Organisation of the paper
  • the Fourier and Mellin transform of $\rho^{(c)}$ in terms of contour integrals (Theorem 2.13 and 2.15).
In Section 3, we will give two contour integral expressions for multivariate Bessel function and the Heckman-Opdam hypergeometric function (Theorem 3.1 and 3.2). Using these and together with Theorem 2.13 and 2.15, we provide proofs for Theorem 1.1 and 1.2. We then end the paper with some Discussions in Section 4.
@@ 2. Dirichlet process and the Markov-Krein correspondence
In this section we review the theory of Dirichlet process based on the notations used in [41, 15]. The purposes of this section are in three folds: to familiarise the readers with the theory of Dirichlet process; to fill in some elementary gaps in some proofs of [41] to make it self-consistent; and to introduce several new theorems (Corollary 2.5, Theorem 2.7,Theorem 2.13 and 2.15).
@@ 2.1. Dirichlet process and its random mean
In general there is no guarantee of the existence of $\rho^{(c)}$, simply because the integral $\int_{\mathbb{R}}x\,D_{c\rho}(\omega)(\mathrm{d}x)$ may not converge for all $D_{c\rho}(\omega)$. In [41, Theorem 1] it is shown that this integral is absolutely integrable almost surely if and only if the logarithm function is integrable with respect to the measure $\rho$, and thus $\rho^{(c)}$ is well-defined.
Theorem 2.1 [41] Existence of the random mean).
@@
[…]
We will briefly illustrate how one shows the existence of $\rho^{(c)}$ in Theorem 2.2; further details can be checked in [41]. The technique used to extend the finitely supported case to the general case is to use an extended Lauricella formula, which has been stated as [41, Eqn. (3.1)].
2.2. The use of the Lauricella formula
Let us firstly consider the simple case when $\rho$ is a discrete measure, finitely supported on $\{x_{1},\ldots,x_{n}\}\subset\mathbb{R}$, given by
$\rho(\mathrm{d}x)=\sum_{j=1}^{n}a_{j}\delta_{x_{j}}(\mathrm{d}x)$ (2.8)
$\int_{\mathbb{R}^{n}}\frac{\mathrm{Dir}_{(ca_{1},\ldots,ca_{n})}(\mathrm{d}\sigma_{1},\ldots,\mathrm{d}\sigma_{n})}{(z-\sum_{j=1}^{n}x_{j}\sigma_{j})^{c}}$ $=\prod_{j=1}^{n}\frac{1}{(z-x_{j})^{a_{j}}}=\exp\left(-c\int_{\mathbb{R}}\log(z-s)\rho(\mathrm{d}s)\right).$ (2.9)
[…]
$\int_{\mathbb{R}^{n}}\varphi\left(\sum_{j=1}^{n}x_{j}\sigma_{j}\right)\mathrm{Dir}_{(ca_{1},\ldots,ca_{n})}(\mathrm{d}\sigma_{1},\ldots,\mathrm{d}\sigma_{n})$ $=\mathbb{E}\,\left[\varphi\left(\int_{\{x_{1},\ldots,x_{n}\}}xD_{c\rho}(\mathrm{d}x)\right)\right]$ (2.11)
holds for any bounded measurable function $\varphi$. Combining (2.9) and (2.11) we have
$\mathbb{E}\,\left[\left(z-\int_{\{x_{1},\ldots,x_{n}\}}xD_{c\rho}(\mathrm{d}x)\right)^{-c}\right]=\exp\left(-c\int_{\mathbb{R}}\log(z-s)\rho(\mathrm{d}s)\right)$ (2.12)
[…]
$\xi_{n}(\mathbb{R})=\left\{y_{j}:j\in J_{n}\right\}$ (2.13)
[…]
$\int_{\mathbb{R}}\varphi(x)\xi_{n,\sharp}\rho(\mathrm{d}x)=\int_{\mathbb{R}}\varphi(\xi_{n}(x))\rho(\mathrm{d}x),$ (2.15)
Lemma 2.3 (Eq. (3.1) in [41]).
$\mathbb{E}\,\left[\left(z-\int_{\mathbb{R}}\varphi(\xi_{n}(x))D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]=\exp\left(-c\int_{\mathbb{R}}\log(z-\varphi(s))\xi_{n,\sharp}\rho(\mathrm{d}s)\right).$ (2.16)
Proof.
$(\sigma_{j})_{j\in J_{n}}=(D_{c\rho}(\xi_{n}^{-1}(y_{j})))_{j\in J_{n}}$ (2.17)
$\int_{\mathbb{R}}\varphi(\xi_{n}(x))D_{c\rho}(\mathrm{d}x)=\sum_{j\in J_{n}}\varphi(y_{j})\sigma_{j}$ (2.18)
The $j$-sum is finite and hence there is no convergence issue. Then by the Lauricella formula (2.9) we have
$\mathbb{E}\,\left[\left(z-\int_{\mathbb{R}}\varphi(\xi_{n}(x))D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]$ $=\int_{\mathbb{R}^{n}}\frac{\mathrm{Dir}_{(ca_{j})_{j\in J_{n}}}(\mathrm{d}\vec{\sigma})}{(z-\sum_{j\in J_{n}}\varphi(y_{j})\sigma_{j})^{c}}=\prod_{j\in J_{n}}\frac{1}{(z-\varphi(y_{j}))^{ca_{j}}}.$ (2.19)
Since $a_{j}=\rho(\xi_{n}^{-1}(y_{j}))$, one has
$\begin{split}\prod_{j\in J_{n}}\frac{1}{(z-\varphi(y_{j}))^{ca_{j}}}&=\exp\left(-c\sum_{j\in J_{n}}\log(z-\varphi(y_{j}))\rho(\xi_{n}^{-1}(y_{j}))\right)\\ &=\exp\left(-c\int_{\mathbb{R}}\log(z-\varphi(\xi_{n}(s)))\rho(\mathrm{d}s)\right)=\exp\left(-c\int_{\mathbb{R}}\log(z-\varphi(s))\xi_{n,\sharp}\rho(\mathrm{d}s)\right).\end{split}$ (2.20)
Combining (2.19) and (2.20) finishes the proof. ∎
2.3. Proof of Theorem 2.1 and Theorem 2.2
Although the proofs are already contained in [41], we attach them here only for the sake of completeness.
Proof of Theorem 2.1.
Using (2.16) and taking $z=1$ and $\varphi(x)=-|x|$, one obtains
$\mathbb{E}\,\left[\left(1+\int_{\mathbb{R}}|\xi_{n}(x)|D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]=\exp\left(-c\int_{\mathbb{R}}\log(1+|s|)\xi_{n,\sharp}\rho(\mathrm{d}s)\right).$ (2.21)
$\mathbb{E}\,\left[\left(1+\lim_{n\to+\infty}\int_{\mathbb{R}}|\xi_{n}(x)|D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]$ $=\exp\left(-c\int_{\mathbb{R}}\log(1+|s|)\rho(\mathrm{d}s)\right).$ (2.22)
On the other hand, assuming $\int_{\mathbb{R}}|x|D_{c\rho}(\mathrm{d}x)$ is finite almost surely, then by a similar argument the right side of (2.22) is non-zero, and thus we have (2.4). ∎
Proof of Theorem 2.2.
$\int_{\mathbb{R}}\frac{\rho^{(c)}(\mathrm{d}x)}{(z-ux)^{c}}=\exp\left(-c\int_{\mathbb{R}}\log(z-us)\rho(\mathrm{d}s)\right)$ (2.23)
For the existence of $\rho^{(c)}$ let us again start with (2.16) and take $\varphi(x)=ux$. Now one takes the limit $n\to+\infty$ on both sides, and again by the dominated convergence theorem we obtain
$\mathbb{E}\,\left[\left(z-u\lim_{n\to+\infty}\int_{\mathbb{R}}\xi_{n}(x)D_{c\rho}(\mathrm{d}x)\right)^{-c}\right]$ $=\exp\left(-c\int_{\mathbb{R}}\log(z-us)\rho(\mathrm{d}s)\right).$ (2.24)
@@ 2.4. Convergence of measures
Theorem 2.1 tells us that $\rho^{(c)}$ is well-defined, namely, the integral $\mathrm{ev}^{\mathrm{id}}(p)$ converges absolutely for any $p\in D_{c\rho}(\Omega)\subset\mathcal{P}(\mathbb{R})$, if and only if (2.4) holds for $\rho$. Thus we have the following commutative diagram
@@
$=\int_{\Omega}\varphi((\mathrm{ev}^{\mathrm{id}}\circ D_{c\rho_{n}})(\omega))\,\mathbb{P}(\mathrm{d}\omega)$
By Theorem 2.1 we know that when $\rho_{n}$ satisfies (2.4), $(\mathrm{ev}^{\mathrm{id}}\circ D_{c\rho_{n}})(\omega)<+\infty$ almost surely, and thus $\varphi(\mathrm{ev}^{\mathrm{id}}(p))$ is almost surely a bounded function in $p$ (the continuity of this function can be verified easily). Then by [15, Thm 1.1] one has
$\lim_{n\to\infty}\int_{\mathcal{P}(\mathbb{R})}\varphi(\mathrm{ev}^{\mathrm{id}}(p))\,(D_{c\rho_{n}})_{\sharp}\mathbb{P}(\mathrm{d}p)=\int_{\mathcal{P}(\mathbb{R})}\varphi(\mathrm{ev}^{\mathrm{id}}(p))\,(D_{c\rho})_{\sharp}\mathbb{P}(\mathrm{d}p),$ (2.27)
@@ 2.5. Further properties of the Markov-Krein correspondence
$\mathcal{V}\,\,:=\left\{\rho:\int\log(1+x^{2})\rho(\mathrm{d}x)<+\infty\right\},$ (2.28)
@@
The following theorem shows a bijective relation between $\mathcal{V}$ and $\mathcal{V}_{c}$. The proof can be found in [41, Theorem 2].
@@
We will show a result relating the tail behaviour of $\rho$ and $\rho^{(c)}$. In short, they have the same tail.
@@
We firstly show this result for the case when $\rho$ is finitely supported, i.e. it is given by $\rho=\sum_{j=1}^{N}a_{j}\delta_{x_{j}}$ for $x_{1},\ldots,x_{n}\in\mathbb{R}$ and $a_{1},\ldots,a_{n}\in[0,1]$ with $a_{1}+\ldots+a_{n}=1$. Substituting (2.30) into (2.11), we see that
@@
$\begin{split}\left|\int_{\mathbb{R}\backslash(-n,n]}\varphi(x)(\rho_{n}(\mathrm{d}x)-\rho(\mathrm{d}x))\right|&=\left|\int_{\mathbb{R}\backslash(-n,n]}\varphi(x)\rho(\mathrm{d}x)\right|\\ &\leq\frac{\int_{\mathbb{R}}h(x)\rho(\mathrm{d}x)}{h(n)}.\end{split}$ (2.38)
@@
$\begin{split}\int_{\mathbb{R}}\varphi(x)\rho^{(c)}(\mathrm{d}x)&=\lim_{R\to\infty}\lim_{n\to\infty}\int_{-R}^{R}\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x)\\ &\leq\lim_{R\to\infty}\limsup_{n\to\infty}\int_{-R}^{R}\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x)\\ &\leq\lim_{R\to\infty}\limsup_{n\to\infty}\int_{\mathbb{R}}\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x)=\limsup_{n\to\infty}\int_{\mathbb{R}}\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x).\end{split}$ (2.40)
@@
$\limsup_{n\to\infty}\int\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x)\leq\frac{1}{c}\lim_{n\to\infty}\int\varphi(x)\rho_{n}(\mathrm{d}x)=\frac{1}{c}\int\varphi(x)\rho(\mathrm{d}x)$ (2.41)
@@
We only need to show the inequalities where the right sides are finite. Since both $e^{a|x|}$ and $|x|^{\alpha}$ are non-negative even convex functions, the inequality 2.30 holds and thus by Theorem 2.7 we have the result. ∎
@@ 2.6. Fourier transform of $\rho^{(c)}$
Fourier transform of $\rho^{(c)}$ has been extensively studied in [41, 53, 15]. We aim at showing a contour integral representation which will be further used in the proofs of our main theorems.
$\rho^{[M]}(\mathrm{d}s):=\rho((-\infty,-M])\delta_{-M}(\mathrm{d}s)+\mathbbm{1}_{(-M,M)}(s)\rho(\mathrm{d}s)+\rho([M,+\infty))\delta_{M}(\mathrm{d}s).$ (2.44)
The following Lemma has been used implicitly in [41, Section 7] without giving a proof. It is a special case of Corollary 2.5 (which is based on results in [15]), while here we provide a alternative elementary proof which does not rely on Corollary 2.5 and results in [15].

### v3

@@ 1.2.5. Organisation of the paper
  • the Fourier and Mellin transform of $\rho^{(c)}$ in terms of contour integrals (Theorem 2.9 and 2.11).
In Section 3, we will give two contour integral expressions for multivariate Bessel function and the Heckman-Opdam hypergeometric function (Proposition 3.1 and Proposition 3.2). Using these and together with Theorem 2.9 and 2.11, we provide proofs for Theorem 1.1 and 1.3. We then end the paper with some Discussions in Section 4.
@@ 2. Dirichlet process and the Markov-Krein correspondence
In this section we review the theory of the Dirichlet process using the notation of [53, 21]. Our aim is to familiarise the reader with the basic framework while also introducing several new results, namely Corollary 2.4, Theorem 2.6, Theorem 2.9, and Theorem 2.11. Some of the statements from [53] will be recalled here for completeness. We note that a few proofs in [53] contain minor gaps (in particular, Theorem 2.9 i.e. [53, Theorem 7]). These issues are purely technical and do not affect the essence of the results, while for the sake of a self-contained exposition, we provide complete proofs and fill in the missing details in the Appendix.
@@ 2.1. Dirichlet process and its random mean
In general there is no guarantee of the existence of $\rho^{(c)}$, simply because the integral $\int_{\mathbb{R}}x\,D_{c\rho}(\omega)(\mathrm{d}x)$ may not converge for all $D_{c\rho}(\omega)$. In [53, Theorem 1] it is shown that this integral is absolutely integrable almost surely if and only if the logarithm function is integrable with respect to the measure $\rho$, and thus $\rho^{(c)}$ is well-defined.
Theorem 2.1 ( [53] Existence of the random mean).
@@
for all $z\in\mathbb{C}\setminus\mathbb{R}$. Also this probability measure $\rho^{(c)}$ is given precisely by the random mean of the Dirichlet process $D_{c\rho}$.
We will briefly illustrate how one shows the existence of $\rho^{(c)}$ in Theorem 2.2; further details can be checked in [53]. The technique used to extend the finitely supported case to the general case is to use an extended Lauricella formula, which has been stated as [53, Eqn. (3.1)]. Detailed proofs of Theorem 2.1 and Theorem 2.2 will be given in Appendix A.1.
@@ 2.2. Convergence of measures
Theorem 2.1 tells us that $\rho^{(c)}$ is well-defined, namely, the integral $\mathrm{ev}^{\mathrm{id}}(p)$ converges absolutely for any $p\in D_{c\rho}(\Omega)\subset\mathcal{P}(\mathbb{R})$, if and only if (2.4) holds for $\rho$. Thus we have the following commutative diagram.
@@
$=\int_{\Omega}\varphi((\mathrm{ev}^{\mathrm{id}}\circ D_{c\rho_{n}})(\omega))\,\mathbb{P}(\mathrm{d}\omega).$
By Theorem 2.1 we know that when $\rho_{n}$ satisfies (2.4), $(\mathrm{ev}^{\mathrm{id}}\circ D_{c\rho_{n}})(\omega)<+\infty$ almost surely, and thus $\varphi(\mathrm{ev}^{\mathrm{id}}(p))$ is almost surely a bounded function in $p$ (the continuity of this function can be verified easily). Then by [21, Thm 1.1] one has
$\lim_{N\to+\infty}\int_{\mathcal{P}(\mathbb{R})}\varphi(\mathrm{ev}^{\mathrm{id}}(p))\,(D_{c\rho_{n}})_{\sharp}\mathbb{P}(\mathrm{d}p)=\int_{\mathcal{P}(\mathbb{R})}\varphi(\mathrm{ev}^{\mathrm{id}}(p))\,(D_{c\rho})_{\sharp}\mathbb{P}(\mathrm{d}p),$ (2.10)
@@ 2.3. Further properties of the Markov-Krein correspondence
$\mathcal{V}\,\,:=\left\{\rho:\int_{\mathbb{R}}\log(1+x^{2})\rho(\mathrm{d}x)<+\infty\right\},$ (2.11)
@@
The following theorem shows a bijective relation between $\mathcal{V}$ and $\mathcal{V}_{c}$. The proof can be found in [53, Theorem 2].
@@
We will show a new result stating that the tail behaviour of $\rho^{(c)}$ is governed by the one for $\rho$.
@@
We firstly show this result for the case when $\rho$ is finitely supported, i.e. it is given by $\rho=\sum_{j=1}^{N}a_{j}\delta_{x_{j}}$ for $x_{1},\ldots,x_{n}\in\mathbb{R}$ and $a_{1},\ldots,a_{n}\in[0,1]$ with $a_{1}+\ldots+a_{n}=1$. Substituting (2.13) into (A.4), we see that
@@
$\begin{split}\left|\int_{\mathbb{R}\setminus(-n,n]}\varphi(x)(\rho_{n}(\mathrm{d}x)-\rho(\mathrm{d}x))\right|&=\left|\int_{\mathbb{R}\setminus(-n,n]}\varphi(x)\rho(\mathrm{d}x)\right|\\ &\leq\frac{\int_{\mathbb{R}}h(x)\rho(\mathrm{d}x)}{h(n)}.\end{split}$ (2.21)
@@
$\begin{split}\int_{\mathbb{R}}\varphi(x)\rho^{(c)}(\mathrm{d}x)&=\lim_{R\to\infty}\lim_{N\to+\infty}\int_{-R}^{R}\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x)\\ &\leq\lim_{R\to\infty}\limsup_{N\to+\infty}\int_{-R}^{R}\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x)\\ &\leq\lim_{R\to\infty}\limsup_{N\to+\infty}\int_{\mathbb{R}}\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x)=\limsup_{N\to+\infty}\int_{\mathbb{R}}\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x).\end{split}$ (2.23)
@@
$\limsup_{N\to+\infty}\int\varphi(x)\rho^{(c)}_{n}(\mathrm{d}x)\leq\frac{1}{c}\lim_{N\to+\infty}\int\varphi(x)\rho_{n}(\mathrm{d}x)=\frac{1}{c}\int\varphi(x)\rho(\mathrm{d}x),$ (2.24)
@@
We only need to show the inequalities where the right sides are finite. Since both $e^{a|x|}$ and $|x|^{\alpha}$ are non-negative even convex functions, the inequality 2.13 holds and thus by Theorem 2.6 we have the result. ∎
@@ 2.4. Fourier transform of $\rho^{(c)}$
Fourier transform of $\rho^{(c)}$ has been extensively studied in [53, 71, 21]. We will show a contour integral representation for the Fourier transform of $\rho^{(c)}$, which will be useful in the proofs of our main theorems.
We say the contour $\mathcal{C}$ is the Hankel loop, if it starts and ends at -$\infty$, enclosing $(-\infty,x]$ oriented counter-clockwise. A typical choice is that, for some $M>x$ it goes horizontally from $\infty-iM$ to MiMM-iM, then vertically to $M+iM$, and then horizontally to $\infty+iM$; see Fig. 1.
[…]
Figure 1. A standard choice for the Hankel loop $\mathcal{C}$.
Also, we denote a sign function for complex numbers as
Theorem 2.9.
Let $u\in\mathbb{C}\setminus\{0\}$. Assuming that
$\int_{\mathbb{R}}e^{x\mathrm{Re}\,u}\,\rho(\mathrm{d}x)<+\infty,$ (2.28)
then
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{xu}=\lim_{M\to+\infty}\frac{\Gamma(c)}{u^{c-1}}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int_{\mathbb{R}}\log(z-s)\rho^{[M]}(\mathrm{d}s)\right)$ (2.29)
$\int_{\mathbb{R}}\rho^{(c)}(\mathrm{d}x)e^{xu}=\frac{\Gamma(c)}{u^{c-1}}\frac{1}{2\pi i}\int_{\mathrm{sgn}(u)\mathcal{C}}\mathrm{d}z\,\exp\left(uz-c\int_{\mathbb{R}}\log(z-s)\rho(\mathrm{d}s)\right).$ (2.30)
Theorem 2.9 might be formally derived from a slightly different form in [53, Theorem 7] using a change of variable, but in the arguments of [53] the convergence of integrals have not been taken care of. We will provide an detailed proof in Appendix A.2, taking care of all issues of convergence.
2.5. Mellin transform of $\rho^{(c)}$
Similar to the Fourier transform, the Mellin transform of $\rho^{(c)}$ can also be rewritten into a contour integral formula, which has not been discussed in the literature.